OpenStudy (anonymous):

How do I simplify 3^-4

OpenStudy (anonymous):

3^-4 = 1/3^4 = 1/(3*3*3*3) = 1/9*9 = 1/81

OpenStudy (anonymous):

Thank you!! Can you please explain how you did it?

OpenStudy (anonymous):

ya sure :) do you know this rule of exponent a^-m = 1/a^m

OpenStudy (anonymous):

nooo, but I do now :)

OpenStudy (anonymous):

good i have just applied that :D

OpenStudy (anonymous):

Ohh thank you!!

OpenStudy (anonymous):

yw xD